# Muffled Murmurs: Environmental effects in the LISA stochastic signal from stellar-mass black hole binaries

### v1

@@ B.2 Validity of gas accretion modeling
[…]
@@
Naively, one might right down the energy balance law as
@@
where $\dot{E}_{\rm GW}$ and $\dot{E}_{\rm Acc}$ are the orbit-averaged outgoing energy fluxes due to GW radiation reaction and accretion-induced drag (that depends linearly on $\xi$, see Eq. 13) respectively. For each term, we have included the corresponding perturbation parameters for book-keeping. Applying Eq. 36 gives
[…]
where we have restored factors of $G$. Relative to the Keplerian dynamics, we have only kept $\mathcal{O}(\epsilon_{\rm RR})$ and $\mathcal{O}(\epsilon_{\rm Acc})$ contributions, and neglected higher order $\mathcal{O}(\epsilon_{\rm Acc}\epsilon_{\rm RR})$, $\mathcal{O}(\epsilon_{\rm RR}^{2})$, and $\mathcal{O}(\epsilon_{\rm Acc}^{2})$ terms because we are effectively performing a leading order expansion in a multiple-timescale analysis [69] (see also [96, 22] for other applications of such an approach). Thus, it is valid to evaluate Eq. 37 with the initial value of the masses. On the other hand, the angular momentum balance law is given by
@@
where $\dot{L}_{\rm GW}$ and $\dot{L}_{\rm Acc}$ are the orbit-averaged outgoing angular momentum fluxes due to GW radiation reaction and accretion-induced drag (that depends linearly on $\xi$, see Eq. 13) respectively. Applying Eq. 38 results in
[…]
@@
which then (for quasi-circular orbits) gives the same back-reaction as when using the angular momentum balance law of Eq. 38, resolving the inconsistency.
We note that a similar problem arises in the case of varying-$G$ theories [51, 90], with the Hamiltonian being time-dependent owing to the time-varying $G(t)$. When using the incorrect energy balance law $\dot{E}_{b}=-\epsilon_{\rm RR}\dot{E}_{\rm GW}$ (the equivalent of Eq. 36 with no drag-induced flux), this leads to an incorrect back-reaction that propagates to the GW phase, which was obtained by [51]. The back-reaction for time-dependent $G(t)$ was subsequently corrected by [90], which is consistent with our discussion for accretion (see also [21]).
@@ C.1 RPL model construction
We note that Eq. 17 is similar to ppE models [101, 47], in the weak coupling limit $f_{s}\gg f_{\rm turn}$ (or $\alpha\ll 1$), corresponding to the GW dominant regime. In this regime, the SGWB model can be used for probing GR deviations in SGWB signals [47]. However, in our case, the energy dissipation due to the environmental is not necessarily small compared to GW emisssion, making the ppE model not applicable for typical astrophysical systems. In fact, the RPL model is effectively a lowest order $[0/1]$ Padé-like approximant [^3] The RPL model is a $[0/1]$ Padé approximant in terms of $f^{1/3}$ and when excluding log-terms in the rational polynomial ansatz., thereby resumming the ppE model of [101, 47] to be valid in the $f_{s}\ll f_{\rm turn}$ regime. Unlike the broken power-law model used by [49], our RPL model is $C^{\infty}$ and smoothly connects the low and high frequency asymptotic regimes.
@@ C.2 RPLP model construction
Specifically, we include an interpolating function that captures the non-perturbative behavior for $f_{s}\sim f_{\rm turn}$, and smoothly vanishes in the $f_{s}\ll f_{\rm turn}$ and $f_{s}\gg f_{\rm turn}$ limits. We accomplish this using the RPLP model that includes a Gaussian correction to the denominator of the RPL model, as given in Eq. 18. The Gaussian correction depends on the parameters $A_{g}(\alpha)$, $f_{\rm peak}(\alpha)$, and $\sigma$. To motivate the functional form of $f_{\mathrm{peak}}(\alpha)$, we compare it with the turning frequency $f_{\mathrm{turn}}(\alpha)$, defined as the frequency at which the energy dissipation due to environmental effects equals that of GW emission, i.e., $|\dot{E}_{\rm env}|=|\dot{E}_{\rm GW}|$ and as shown in Eqs. 11 and 16. The quantity $f_{\mathrm{turn}}(\alpha)$ is obtained by numerical evaluation, redshifted from the source frame to the observer frame by dividing by $(1+z)$, and averaged over the population. We find that the fitted $f_{\mathrm{peak}}(\alpha)$ and $f_{\mathrm{turn}}(\alpha)$ are both expressed as power laws in $\alpha$, and that they share the same exponent, with amplitudes differing only by a factor $\sim 2$. In addition, we fit the amplitude $A_{g}(\alpha)$ and determine $\sigma$ using the full width at half maximum (FWHM) method. We find that $\sigma$ remains approximately constant across different values of $\alpha$.
@@
For the cases of dynamical friction and gas accretion, we show a comparison of both the RPL and RPLP models against the true signal in Fig. 11. Recall that the true signal for dynamical friction and gas accretion are generated using methods described in Sec. III.1 and Sec. III.2 respectively. We fiducially set $\rho=10^{-7}\mathrm{g}\,\mathrm{cm}^{-3}$ for dynamical friction and $f_{\rm Edd}=10$ for gas accretion. The RPL model (blue dashed line) has a significant mismatch with the true signal (black solid line) in the intermediate regime, for both dynamical friction and gas accretion. However, the RPLP model (dashed red line) effectively removes this mismatch at intermediate frequencies.
In the case of dynamical friction, a small mismatch arises at low frequencies due to neglecting the $(1+z)$ factor in $\ln[(1+z)f]$, as shown in Fig. 11. We do this simplification to scale out the frequency dependence when computing the asymptotic value of $A_{m}$, and thus underestimating the signal in the low frequency regime. Note that for gas accretion, the RPLP model matches the signal well in the low frequency asymptotic regime since there are no $\ln[(1+z)f]$ terms involved. Thus for gas accretion, the frequency dependence can be trivially scaled out, resulting in a highly accurate asymptotic estimate of $A_{m}$.
@@ C.3 Bayesian parameter estimation checks with the RPLP model
For the case of dynamical friction, since the RPLP model underestimates the true signal at low frequencies, we incurred biases in the marginalized posterior of $\rho$ particularly for large $\rho^{\rm inj}$, as shown in Fig. 3. Notwithstanding these model induced errors, the posterior predictive of the RPLP model agrees reasonably well with the true signal, as we show in Fig. 12.
@@
In Fig. 3, we had also found that the marginalized posterior of $A_{\rm vac}$ is consistently biased to smaller values with increasing $\rho^{\rm inj}$. To isolate this behavior from the bias incurred due to mismatch with the true signal, we performed an injection–recovery check with the RPLP model. In Fig. 13, we show the marginalized posteriors of $A_{\rm vac}$ and $\rho$ with the Galactic foreground parameters fixed. We no longer see any bias in the recovery of $\rho$, implying that indeed the bias we found in Fig. 3 is mainly due to the mismatch between the RPLP model and the signal (as illustrated in Fig. 11).
Figure 13: Marginalized posterior probability of $A_{\rm vac}$ and $\rho$ with mock data generated using the RPLP model. The posteriors for different injected densities are obtained with Galactic foreground parameters fixed. Two-dimensional contours correspond to $90\%$ credible regions. Dash-dotted lines represent the injected value for $\rho$ and $A_{\rm vac}$. Observe that the density is better recovered compared to Fig. 3.
We observe that the bias in the marginalized posterior of $A_{\rm vac}$ (relative to its asymptotic vacuum value) in Fig. 13 is smaller than in Fig. 3. Yet, there is still a significant bias in $A_{\rm vac}$ for $\rho^{\rm inj}=10^{-7}\mathrm{g}\,\mathrm{cm}^{-3}$. This bias cannot be simply explained by the failure of the RPL model, since the injection here is with the same model and we do not observe significant biases with $\rho$. We can however explain the $A_{\rm vac}$ bias in the following way. The asymptotic vacuum value for $A_{\rm vac}$ is meaningful provided that the high frequency asymptotic regime $f_{s}\gg f_{\rm turn}$ is within the LISA band. For $\rho^{\rm inj}=10^{-7}\mathrm{g}\,\mathrm{cm}^{-3}$, $f_{\rm turn}$ shifts to the higher frequencies, thereby pushing the high frequency asymptotic regime to the less sensitive part of the LISA band. Therefore, using the asymptotic value of $A_{\rm vac}$ as truth to assess bias is not completely valid. Thus, this contributes to the biases observed in Fig. 3 for $A_{\rm vac}$. Put another way, for such large $\rho^{\rm inj}$ the validity regime of the asymptotic matching that is built into the RPLP model affects the inference.

### v2

@@ B.2 Validity of gas accretion modeling
$\begin{aligned} f_{\rm Edd}&\lesssim\left(1.4\times 10^{7}\right)\left(\dfrac{f_{s}}{10^{-5}\,\mathrm{Hz}}\right)\left(\dfrac{\tau}{4.5\times 10^{7}\,\mathrm{yr}}\right)\left(\dfrac{8}{5+3\xi}\right).\end{aligned}$ (35)
Thus for astrophysical values of $f_{\rm Edd}\sim 0.01$$100$, the effects of accretion are well within the perturbative regime.
@@
Naively, one might write down the energy balance law as
@@
where $\dot{E}_{\rm GW}$ and $\dot{E}_{\rm Acc}$ are the orbit-averaged outgoing energy fluxes due to GW radiation reaction and accretion-induced drag (that depends linearly on $\xi$, see Eq.˜13) respectively. For each term, we have included the corresponding perturbation parameters for book-keeping. Applying Eq.˜36 gives
$\begin{aligned} \left(\dfrac{df_{s}}{dt}\right)_{E}&=\epsilon_{\rm RR}\dfrac{96}{5}\pi^{8/3}G^{5/3}\mathcal{M}_{0}^{5/3}f_{s}^{11/3}\\ &+\epsilon_{\rm Acc}\dfrac{(-5+6\xi)}{2}\dfrac{f_{\rm Edd}}{\tau}f_{s},\end{aligned}$ (37)
where we have restored factors of $G$. Relative to the Keplerian dynamics, we have only kept $\mathcal{O}(\epsilon_{\rm RR})$ and $\mathcal{O}(\epsilon_{\rm Acc})$ contributions, and neglected higher order $\mathcal{O}(\epsilon_{\rm Acc}\epsilon_{\rm RR})$, $\mathcal{O}(\epsilon_{\rm RR}^{2})$, and $\mathcal{O}(\epsilon_{\rm Acc}^{2})$ terms because we are effectively performing a leading order expansion in a multiple-timescale analysis [76] (see also [104, 23] for other applications of such an approach). Thus, it is valid to evaluate Eq.˜37 with the initial value of the masses. On the other hand, the angular momentum balance law is given by
@@
where $\dot{L}_{\rm GW}$ and $\dot{L}_{\rm Acc}$ are the orbit-averaged outgoing angular momentum fluxes due to GW radiation reaction and accretion-induced drag (that depends linearly on $\xi$, see Eq.˜13) respectively. Applying Eq.˜38 results in
$\begin{aligned} \left(\dfrac{df_{s}}{dt}\right)_{L}&=\epsilon_{\rm RR}\dfrac{96}{5}\pi^{8/3}G^{5/3}\mathcal{M}_{0}^{5/3}f_{s}^{11/3}\\ &+\epsilon_{\rm Acc}(5+3\xi)\dfrac{f_{\rm Edd}}{\tau}f_{s},\end{aligned}$ (39)
@@
which then (for quasi-circular orbits) gives the same back-reaction as when using the angular momentum balance law of Eq.˜38, resolving the inconsistency.
We note that a similar problem arises in the case of varying-$G$ theories [59, 98], with the Hamiltonian being time-dependent owing to the time-varying $G(t)$. When using the incorrect energy balance law $\dot{E}_{b}=-\epsilon_{\rm RR}\dot{E}_{\rm GW}$ (the equivalent of Eq.˜36 with no drag-induced flux), this leads to an incorrect back-reaction that propagates to the GW phase, which was obtained by [59]. The back-reaction for time-dependent $G(t)$ was subsequently corrected by [98], which is consistent with our discussion for accretion (see also [22]).
@@ C.1 RPL model construction
We note that Eq.˜17 is similar to ppE models [109, 55], in the weak coupling limit $f_{s}\gg f_{\rm turn}$ (or $\alpha\ll 1$), corresponding to the GW dominant regime. In this regime, the SGWB model can be used for probing GR deviations in SGWB signals [55]. However, in our case, the energy dissipation due to the environmental is not necessarily small compared to GW emisssion, making the ppE model not applicable for typical astrophysical systems. In fact, the RPL model is effectively a lowest order $[0/1]$ Padé-like approximant [^4] The RPL model is a $[0/1]$ Padé approximant in terms of $f^{1/3}$ and when excluding log-terms in the rational polynomial ansatz., thereby resumming the ppE model of [109, 55] to be valid in the $f_{s}\ll f_{\rm turn}$ regime. Unlike the broken power-law model used by [57], our RPL model is $C^{\infty}$ and smoothly connects the low and high frequency asymptotic regimes.
@@ C.2 RPLP model construction
Specifically, we include an interpolating function that captures the non-perturbative behavior for $f_{s}\sim f_{\rm turn}$, and smoothly vanishes in the $f_{s}\ll f_{\rm turn}$ and $f_{s}\gg f_{\rm turn}$ limits. We accomplish this using the RPLP model that includes a Gaussian correction to the denominator of the RPL model, as given in Eq.˜18. The Gaussian correction depends on the parameters $A_{g}(\alpha)$, $f_{\rm peak}(\alpha)$, and $\sigma$. To motivate the functional form of $f_{\mathrm{peak}}(\alpha)$, we compare it with the turning frequency $f_{\mathrm{turn}}(\alpha)$, defined as the frequency at which the energy dissipation due to environmental effects equals that of GW emission, i.e., $|\dot{E}_{\rm env}|=|\dot{E}_{\rm GW}|$ and as shown in Eqs.˜11 and 16. The quantity $f_{\mathrm{turn}}(\alpha)$ is obtained by numerical evaluation, redshifted from the source frame to the observer frame by dividing by $(1+z)$, and averaged over the population. We find that the fitted $f_{\mathrm{peak}}(\alpha)$ and $f_{\mathrm{turn}}(\alpha)$ are both expressed as power laws in $\alpha$, and that they share the same exponent, with amplitudes differing only by a factor $\sim 2$. In addition, we fit the amplitude $A_{g}(\alpha)$ and determine $\sigma$ using the full width at half maximum (FWHM) method. We find that $\sigma$ remains approximately constant across different values of $\alpha$.
@@
For the cases of dynamical friction and gas accretion, we show a comparison of both the RPL and RPLP models against the true signal in Fig.˜11. Recall that the true signal for dynamical friction and gas accretion are generated using methods described in Sec. III.1 and Sec. III.2 respectively. We fiducially set $\rho=10^{-7}\mathrm{g}\,\mathrm{cm}^{-3}$ for dynamical friction and $f_{\rm Edd}=10$ for gas accretion. The RPL model (blue dashed line) has a significant mismatch with the true signal (black solid line) in the intermediate regime, for both dynamical friction and gas accretion. However, the RPLP model (dashed red line) effectively removes this mismatch at intermediate frequencies.
In the case of dynamical friction, a small mismatch arises at low frequencies due to neglecting the $(1+z)$ factor in $\ln[(1+z)f]$, as shown in Fig.˜11. We do this simplification to scale out the frequency dependence when computing the asymptotic value of $A_{m}$, and thus underestimating the signal in the low frequency regime. Note that for gas accretion, the RPLP model matches the signal well in the low frequency asymptotic regime since there are no $\ln[(1+z)f]$ terms involved. Thus for gas accretion, the frequency dependence can be trivially scaled out, resulting in a highly accurate asymptotic estimate of $A_{m}$.
@@ C.3 Bayesian parameter estimation checks with the RPLP model
For the case of dynamical friction, since the RPLP model underestimates the true signal at low frequencies, we incurred biases in the marginalized posterior of $\rho$ particularly for large $\rho^{\rm inj}$, as shown in Fig.˜3. Notwithstanding these model induced errors, the posterior predictive of the RPLP model agrees reasonably well with the true signal, as we show in Fig.˜12.
@@
In Fig.˜3, we had also found that the marginalized posterior of $A_{\rm vac}$ is consistently biased to smaller values with increasing $\rho^{\rm inj}$. To isolate this behavior from the bias incurred due to mismatch with the true signal, we performed an injection–recovery check with the RPLP model. In Fig.˜13, we show the marginalized posteriors of $A_{\rm vac}$ and $\rho$ with the Galactic foreground parameters fixed. We no longer see any bias in the recovery of $\rho$, implying that indeed the bias we found in Fig.˜3 is mainly due to the mismatch between the RPLP model and the signal (as illustrated in Fig.˜11).
Figure 13: Marginalized posterior probability of $A_{\rm vac}$ and $\rho$ with mock data generated using the RPLP model. The posteriors for different injected densities are obtained with Galactic foreground parameters fixed. Two-dimensional contours correspond to $90\%$ credible regions. Dash-dotted lines represent the injected value for $\rho$ and $A_{\rm vac}$. Observe that the density is better recovered compared to Fig.˜3.
We observe that the bias in the marginalized posterior of $A_{\rm vac}$ (relative to its asymptotic vacuum value) in Fig.˜13 is smaller than in Fig.˜3. Yet, there is still a significant bias in $A_{\rm vac}$ for $\rho^{\rm inj}=10^{-7}\mathrm{g}\,\mathrm{cm}^{-3}$. This bias cannot be simply explained by the failure of the RPLP model, since the injection here is with the same model and we do not observe significant biases with $\rho$. We can however explain the $A_{\rm vac}$ bias in the following way. The asymptotic vacuum value for $A_{\rm vac}$ is meaningful provided that the high frequency asymptotic regime $f_{s}\gg f_{\rm turn}$ is within the LISA band. For $\rho^{\rm inj}=10^{-7}\mathrm{g}\,\mathrm{cm}^{-3}$, $f_{\rm turn}$ shifts to the higher frequencies, thereby pushing the high frequency asymptotic regime to the less sensitive part of the LISA band. Therefore, using the asymptotic value of $A_{\rm vac}$ as truth to assess bias is not completely valid. Thus, this contributes to the biases observed in Fig.˜3 for $A_{\rm vac}$. Put another way, for such large $\rho^{\rm inj}$ the validity regime of the asymptotic matching that is built into the RPLP model affects the inference.